A long solenoid has a length of 0.66 m and contains 1500 turns of wire. There is a current of 3.9 A in the wire. What is the magnitude of the magnetic field within the solenoid? Units the tolerance is +/-3% Click if you would like to Show Work for this question: Open Show Work
Answer #1

We know that the magnetic field due to solenoid

B=u° N I /L

here u° =4π ×10^(-7) ,N=1500 , L=0.66 m , I=3.9 A

B=4π×10^(-7)×1500×3.9/0.66

B=111383.74×10^(-7)

B=1.114×10^(-2) T Answer